A train is crossing an observer standing on the platform. The first compartment of the train takes two seconds to cross the observer while the 2nd takes 2.5 seconds to cross him/


train is moving with uniform acceleration. FIND


a) velocity of the train when the front of 1st compartment crosses the observer.


b) acceleration of train


GIVEN THAT LENGHT OF EACH COMPARTMENT IS 15m.

Grade:

1 Answers

Sadhna Agrawal
6 years ago
For the first compartment ;S=15mT=2 sS=ut+1/2at^215=2u+2a eq-1V=u+atV=u+2tFor second compartmentU= V of first compartment =u+2tT=2.5 sS=15mS=ut+1/2at^215=(u+2t)2.5 + 1/2a6.25 eq-2Solving eq.1 and 2 we get;u=8.17 m/sa= -0.67 m/s^2